I tried to solve this problem in many ways, please help
problem (2.2^2)^3 =
my answer 2.2=4
(4^2)=16
16^3=4096
then
(2.2^2)^3
(4.2)
(8)^3
512

: I tried to solve this problem in many ways, please help
problem (2.2^2)^3 =
my answer 2.2=4
(4^2)=16
16^3=4096
then
(2.2^2)^3
(4.2)
(8)^3
512


Answer by checkley71(8405)

(2.2^2)^3
(2.2^2)=4.84
(4.84^3)=113.3799 answer.